### v1

# SECURA: Sigmoid-Enhanced CUR Decomposition with Uninterrupted Retention and Low-Rank Adaptation in Large Language Models

Abstract
With the rapid development of large language models (LLMs), fully fine-tuning(FT) these models has become less feasible due to high computational demands. Moreover, FT also increases the risk of catastrophic forgetting. As an alternative, Low-Rank Adaptation (LoRA) has been proposed. By fine-tuning only a small subset of parameters, LoRA achieves performance similar to FT while significantly reducing resource requirements. However, since LoRA inherits FT’s design, the issue of catastrophic forgetting still remains. To address these limitations, we propose SECURA: Sigmoid-Enhanced CUR Decomposition LoRA, a novel PEFT variant designed to mitigate catastrophic forgetting while improving fine-tuning performance. Our method introduces a new normalization technique, SigNorm, to enhance parameter retention and improve overall performance. SECURA has been evaluated on a diverse range of tasks, including mathematical problem-solving (GSM8K), challenging question-answering (CNNDM), translation (NewsDE), and complex multiple-choice reasoning (LogiQA). Experimental results demonstrate it achieves an average fine-tuning improvement of 3.59% across 4 MCQ sphere tasks and a 2.51% improvement in 5 QA sphere tasks across Gemma2 2b, Qwen2 1.5b, Qwen 2 7b,Llama3 8b,Llama3.1 8b compared to DoRA. Additionally, SECURA demonstrates superior knowledge retention capabilities, achieves state-of-the-art in 16 continual learning tests keep more than 70% accuracy on LLMs basic knowledge compared with Experience Replay(ER), sequential learning(SEQ),EWC, I-LoRA and CUR-LoRA.
SECURA: Sigmoid-Enhanced CUR Decomposition with Uninterrupted Retention and Low-Rank Adaptation in Large Language Models
1 Introduction
With the rapid development of transformer-based generative models, parameter scales have grown exponentially to enhance multi-task capabilities. Notable examples in natural language models(NLP) like LLaMA 3.1 (Dubey et al., 2024) and Qwen 2.5 (Team, 2024) achieve remarkable performance on complex tasks including text generation, logical reasoning, and multilingual processing. However, their massive parameterization (e.g., 70B+ parameters) creates deployment barriers: Full fine-tuning requires over 1.5TB of GPU memory and weeks of computation. Even with distilled versions, they remain prohibitively expensive for most individual developers and small research labs.
Parameter-Efficient Fine-Tuning (PEFT) (Xu et al., 2023) methods, such as Low-Rank Adaptation (LoRA) (Whitehouse et al., 2024), address this challenge. LoRA inject trainable tiny low-rank matrices (typically <0.1% of original parameters) and frozen basic weights to reduce the computational cost of fine-tuning. Performed well in multi-task learning, multilingual summarisation, and transfer learning (Whitehouse et al., 2024; Zhao et al., 2024).
However, LoRA remains an approximation of FT, meaning it hardly match or exceed the performance of full fine-tuning. Furthermore, it also inherits FT’s critical issue: catastrophic forgetting (McCloskey and Cohen, 1989) . While freezing the base model weights alleviates this problem to some extent, the addition of LoRA matrices reintroduces the issue. Additionally, fine-tuning fewer parameters, while reducing computational cost, increases the risk of catastrophic forgetting before the model finds an optimal solution. Although Weight normalization techniques (Salimans and Kingma, 2016) can be used to mitigate these issues but have not provided a comprehensive solution. Existing LoRA variants primarily focus on improving specific aspects: enhancing performance (e.g. DoRA (Liu et al., 2024), AdaLoRA (Zhang et al., 2023)), accelerating convergence (e.g., PiSSA (Meng et al., 2024)), or only reducing catastrophic forgetting (e.g. I-LoRA (Ren et al., 2024), CUR-LoRA (Fawi, 2024)). However, few approaches simultaneously address uninterrupted retention of knowledge and keep the performance.
Motivated by challenges in catastrophic forgetting, we drew inspiration from QKV pruning theory (Lv et al., 2024), and Elastic Weight Consolidation(EWC) (Aich, 2021) to design a new LoRA variant: SECURA. Our approach leverages two core innovations:
  • CABR Decomposition: To enhance the low performance of CUR-LoRA, we introduced an inverse low-rank adaptation matrix into the trainable $U$-matrix. By incorporating additional dimensions, CABR improves CUR-LoRA’s performance, achieving results comparable to or exceeding those of standard LoRA.
  • SigNorm Normalization: To balance model stability and performance, we designed SigNorm—a normalization method using Sigmoid’s gradual transition property. It dynamically adjusts parameters during training to prevent catastrophic forgetting by prioritizing the collective impact of central parameter changes over extreme values, eliminating the need for historical data or prior outputs. This ensures balanced weight updates while maintaining adaptability.
By integrating SigNorm and CABR decomposition, SECURA effectively addresses memory retention challenges in continual learning while surpassing existing LoRA methods and their variants in accuracy under comparable computational budgets. Experiments on 5 LLMs (Gemma2 2B, Qwen2 1.5B/7B, Llama3 8B, Llama3.1 8B) demonstrate SECURA’s superiority: it achieves 3.63% average improvement on 4 multiple-choice (MCQ) tasks and 2.56% on 5 question-answering (QA) tasks over baseline LoRA, even rivaling advanced variants like DoRA. For knowledge retention, SECURA maintains over 70% accuracy on fine-tuned test sets, outperforming state-of-the-art methods such as Experience Replay (ER) (Fedus et al., 2020), sequential learning (SEQ) (Sutskever et al., 2014), EWC (Aich, 2021),I-LoRA (Ren et al., 2024) and CUR-LoRA (Fawi, 2024). These results highlight SECURA’s dual strengths in task adaptation and preserving pre-trained knowledge.
2 Background
Figure 1: SigNorm Update method: The figure illustrates the process flow of the SigNorm Normalization algorithm. Shows the steps of starting with the fusion of the former weight matrix with the CABR module and moving through the ratio loss matrix calculation, the normalization and sigmoid steps, followed by the final limiting of the ratio loss matrix.
Low-Rank Adaption.
LoRA (Whitehouse et al., 2024) is a method that uses low-rank matrices, hypothesizing that in the parameters matrix, only few parameters controls main knowledge. By introducing low-rank parameter matrices A and B, “it can generate a matrix that provides the same output, where A initialized by random sampling, B initialized by zeros.
Formally, consider a weight matrix $\boldsymbol{W}\in\mathbb{R}^{h\times d}$ within the original LLMs. LoRA introduces two low-rank matrices, $\boldsymbol{A}\in\mathbb{R}^{h\times r}$ and $\boldsymbol{B}\in\mathbb{R}^{r\times d}$, where $r\ll\min(h,d)$. Instead of directly updating the weight matrix, LoRA modifies the model’s forward pass according to the following equation:
$\boldsymbol{W}^{\prime}=\boldsymbol{o}+\Delta\boldsymbol{o}=\boldsymbol{W}+\boldsymbol{AB}$
By freezing the basic weight and adding the AB matrix the model can be fine-tuned similarly like FT.
CUR-Decomposition and Its Application in CUR-LoRA.
CUR-Decomposition (Hamm and Huang, 2019) is a matrix factorization method that selects specific rows and columns from a matrix. It is used to decompose a matrix in various ways depending on the design. When combined with gradient descent, it can be applied easily using the following formula:
$\boldsymbol{W}=\boldsymbol{C}\boldsymbol{U}\boldsymbol{R}(\boldsymbol{x}),$
where $\boldsymbol{W}\in\mathbb{R}^{h\times d}$ is the original parameter weight, $\boldsymbol{C}\in\mathbb{R}^{h\times r}$ and $\boldsymbol{R}\in\mathbb{R}^{r\times d}$ are the selected columns and rows from $\boldsymbol{W}$, and $\boldsymbol{U}\in\mathbb{R}^{r\times r}$ is a learnable matrix. Here, $r\ll\min(h,d)$.
CUR-LoRA (Fawi, 2024) builds on CUR-Decomposition, by selecting the least important rows and columns as $\boldsymbol{C}$ and $\boldsymbol{R}$. These selected rows and columns are much fewer than the original matrix and have the lowest norms, which leads to slower updates. This decomposition allows CUR-LoRA to focus on updating the less important parts of the matrix, because less important parameters are more sensitive to slower updates, while important ones are less affected. In this way, the model can preserve the previously learned knowledge while efficiently learning new information, improving both training efficiency and performance.
Dynamic Network Pruning.
Dynamic Network Pruning (Lin et al., 2020) is an input-adaptive method that improves the efficiency of neural networks by dynamically adjusting their structure during inference. Unlike static pruning (Siciliano et al., 2024), which removes redundant weights or neurons, dynamic pruning adapts the network based on the characteristics of each input. This approach allows the network to selectively activate relevant parameters, optimizing computational cost.
Formally, consider a weight matrix $\boldsymbol{W}\in\mathbb{R}^{h\times d}$ and a binary mask matrix $\boldsymbol{M}(\boldsymbol{x})\in\{0,1\}^{h\times d}$, where $\boldsymbol{x}$ represents the input data. Dynamic pruning modifies the forward pass as follows:
$\boldsymbol{y}=\sigma\left((\boldsymbol{W}\odot\boldsymbol{M}(\boldsymbol{x}))\boldsymbol{x}\right)$
where $\odot$ denotes element-wise multiplication, and $\sigma$ represents the activation function.
The mask $\boldsymbol{M}(\boldsymbol{x})$ is dynamically generated based on input-dependent criteria such as weight importance, gradient magnitudes, or gating mechanisms. This ensures that only a subset of parameters is used for each input.
Dynamic pruning offers significant computational efficiency by reducing unnecessary operations, allowing the network to adjust its complexity based on input difficulty. However, generating $\boldsymbol{M}(\boldsymbol{x})$ and ensuring efficient real-time pruning remain active research challenges.
3 Methodology
Our proposed SECURA method is based on three key components: CABR-LoRA reverse decomposition fine-tuning, the SigNorm Normalization algorithm, and two types of SECURA Merge method. Drawing inspiration from dynamic network pruning and CUR-decomposition, we hypothesize that parameters with higher weights in the original model are more important, as they store more critical knowledge. By constraining these important parameters and focusing on fine-tuning the less important ones, SECURA can achieve remarkable performance while significantly mitigating catastrophic forgetting. Appendix B shows all intuitive equation derivation.
Figure 2: Gradient Analysis: The gradient change during training shows LoRA has higher gradient fluctuations, indicating greater parameter reshaping, which may increase the risk of catastrophic forgetting. In contrast, SECURA shows lower gradient changes, avoided excessive drift, suggesting more stable parameter updates, improving catastrophic forgetting mitigation and ability of finding a better optimum for the model’s parameters.
3.1 CABR-LoRA Reverse Decomposition Initialization
As shown in Figure 1, we initialized CABR-LoRA as a reverse LoRA decomposition of the $U$ matrix. This allows the "not important" columns and rows to remain smaller while maintaining a higher-dimensional space for fine-tuning SECURA. The formula for initialize of SECURA is:
$\mathcal{W}_{\text{CABRWeight}}=C\cdot A_{\text{initial}}\cdot B_{\text{zeros}}\cdot R.$ (1)
CABR-LoRA is based on CUR-LoRA, utilizing CUR decomposition to break down the weight matrices. CUR-LoRA reduces the number of trainable parameters to locate lest important $\boldsymbol{C}$ and $\boldsymbol{R}$, but this often leads to the loss of important parameters, causing performance degradation. To address this, we introduce CABR decomposition, which decomposes the $U$ matrix, into two matrices: $\mathcal{W}_{\text{A}}$ and $\mathcal{W}_{\text{B}}$, which shapes are $r\times m$ and $m\times r$. Respectively here $m>r$. $\mathcal{W}_{\text{A}}$ is initialized via SVD decomposition, while $\mathcal{W}_{\text{B}}$ is initialized with zeros. The decomposition follows the formulas below:
$\mathcal{W}_{\text{frozen}}=USV^{T}.$ (2)
$\mathcal{W}_{A}=\mathcal{U}_{\text{[:r,:]}}\cdot\mathcal{S}_{\text{[:,:]}}\cdot V_{\text{[:m,:]}}^{T}$ (3)
$\mathcal{W}_{B}=\mathcal{M}_{\text{Zeros}},\quad\text{where}\quad\mathcal{M}_{\text{Zeros}}\in\mathbb{R}^{m\times r}.$ (4)
By initialize $\mathcal{W}_{A}$ with SVD decomposition, CABR-LoRA preserved the most knowledge learned before. The initialization of $\mathcal{W}_{B}$ with zeros ensures that this matrix will fine-tune from the base LLM without altering the critical knowledge.
3.2 SigNorm Normalization Algorithm
The CABR decomposition provides a higher dimension of fine-tuning, but only using CABR-LoRA alone does not fully address catastrophic forgetting, which is crucial for performance in lifelong learning. To tackle this, we introduce the SigNorm Normalization algorithm, inspired by dynamic network pruning. This algorithm adjusts the weight updates by controlling the change magnitude between the former and new weights, provide a stable parameter update figure 2.
The core idea is to apply a Sigmoid function to dynamically adjust the magnitude of weight updates by getting a weight pruning matrix which value between range in $[1,2]$. Because sigmoid function mainly focus on central value’s change and suppresses outliers on both ends. We can allow SECURA deal with whole matrix’s changing tendency while limiting the extreme parameters’ effect. At the same time, restricting the update amplitude of lower-norm columns and rows ensures that important parameters(with higher norms)’ magnitude experience minimal change, while less important parameters (with smaller norms) undergo more significant updates. The former weight matrix is denoted as $W_{\text{former}}$, and the SECURA’s matrices as $A$ and $B$.
First, the merged weight matrix is computed as:
$\mathcal{W}_{\text{SECURAMerged}}=CABR+\mathcal{W}_{\text{former}}.$ (5)
Next, we calculate the MagLossMatrix $\mathcal{M}_{\text{Mag}}$, which captures the relative magnitude change between the new and former weights. Here $\epsilon$ is a small value to avoid division by too small parameters:
$\mathcal{M}_{\text{Mag}}=\left|\frac{\mathcal{W}_{\text{SECURAMerged}}}{\mathcal{W}_{\text{former}}+\epsilon}\right|.$ (6)
Then, we normalize the MagLossMatrix to the range $[-0.5,0.5]$ to balance the values on both sides of zero:
$\mathcal{M}_{\text{norm}}=\left[\frac{\mathcal{M}_{\text{Mag}}}{\max(\mathcal{M}_{\text{Mag}})+\epsilon}-0.5\right]\cdot\text{Scale}.$ (7)
Note that the Scale parameter here resizes the normed value to ensure it falls within the range $[0,1]$ after applying the Sigmoid function, as the Sigmoid function only produces a range of $[0.3775,0.6225]$ for the $[-0.5,0.5]$ interval.
Finally, we apply the Sigmoid function to the normalized matrix to produce the SigNorm values and get the final restring matrix $\mathcal{M}_{\text{Res}}$:
$\mathcal{M}_{\text{Res}}=2-\sigma\left(\mathcal{M}_{\text{norm}}\right).$ (8)
Once we have weight pruning matrix $\mathcal{M}_{\text{Res}}$ the megred matrix $W_{\text{SECURAMerged}}$, will adjust the update process by dividing the SigNorm pruning-weight $\mathcal{M}_{\text{Res}}$:
$\mathcal{W}_{\text{updated}}=\frac{\mathcal{W}_{\text{SECURAMerged}}}{\mathcal{M}_{\text{Res}}}.$ (9)
Figure  3 shows, this dynamic update process allows controlled pruning of weights, where updates for important parameters are slowed down and less important one remains, helping preserve important knowledge and mitigating catastrophic forgetting.
3.3 Merge with Basic Model:
Limiting parameter growth alone is not the optimal solution. In our early experiments, we found that using only the methods from Sections 3.1 and 3.2 led to worse performance—not only in basic performance but also in mitigating catastrophic forgetting. This issue was identified through insights from delayed updates (Kosta et al., 2020) and Target Network (Gao et al., 2017). We realized that while basic parameters should not remain static during long-term training, they must be updated to prevent performance from stagnating or even deteriorating.
If the basic parameters are kept unchanged, two cycle emerges: initially, unimportant parameters are updated, while important ones are constrained. However, in the second cycle, important parameters become smaller than unimportant ones, and the cycle worsens as important parameters begin to change more rapidly, causing further catastrophic forgetting and worse performance.
To address this issue, we propose two merging methods:
Merge with basic weight(M1 method):
This approach requires only one SECURA training cost but does not keep the basic LLM weights unchanged. The update rule is as follows:
$\mathcal{W}_{\text{Newformer}}=CABR+\mathcal{W}_{\text{former}}.$ (10)
Merge with SECURA(M2 method):
This method uses two identical SECURA models but freezes the base LLM weights. An additional calculation is performed to obtain the final updated weights The update rule is as follows:
$\mathcal{W}_{\text{former}}=CA_{\text{former}}B_{\text{former}}R+\mathcal{W}_{\text{basic}}.$ (11)
The merge process is described by the following formulas:
$A_{\text{Newformer}}=A_{\text{train}}.$ (12)
$B_{\text{Newformer}}=B_{\text{former}}+B_{\text{train}}.$ (13)
In both M1 and M2, the new parameters of $B_{\text{former}}$ must be reset to zero after each update, while the matrix remains unchanged. This helps the LLM learn new knowledge effectively.
M1 offers better fine-tuning but sacrifices some knowledge retention, while M2 retains more knowledge but sacrifices fine-tuning performance due to shallower merging. Based on our ablation study 4, for tasks focused on knowledge retention, we recommend M2, and for tasks prioritizing fine-tuning performance, M1 is preferred. M1 also provides a balance, maintaining a high level of knowledge retention while delivering better fine-tuning results.
Figure 3: Comparison of Weight SVD Norm across different methods: The "Basic" method represents the original LLM weights without fine-tuning. After 2000 training steps, DoRA shows a 0.517 decrease in norm, indicating some loss of information. SECURA preserves more knowledge with a minor 0.061 change, while LoRA shows a significant 3.076 increase, reflecting greater weight modification.
4 Experiments
Figure 4: The performance of SECURA compared to 16 tasks baselines sequentially trained using LLaMA-3 8B (left) and Qwen-2 7B (right) backbones, tested under F-task with learning rate 1e-5. Detailed results are in Appendix E
4.1 Experimental Setup
Datasets.
We conducted comprehensive evaluations on 18 datasets: 15 from BBM (Srivastava et al., 2023), while remaining datasets are selected from Llama3’s (Dubey et al., 2024), Qwen2’s (Team, 2024), Gemma2’s (Team et al., 2024) base training or fine-tuning datasets. These datasets include 8 multiple-choice question (MCQ) datasets and 10 question-answering (QA) datasets, spanning six domains: mathematical reasoning(GSM8K) (Cobbe et al., 2021), logical deduction(LogicalQA) (Srivastava et al., 2023), multilingual identification(CNN/DailyMail) (See et al., 2017), programming challenges(Multipl-E) (Cassano et al., 2022), translation questions(News Commentary dataset) (Tiedemann, 2012), and commonsense understanding(CommonsenseQA) (Talmor et al., 2019). Specifically, all of our benchmarks are carefully handled, with 1000 samples taken from each respective source and a 9:1 training-to-testing split.
To evaluate catastrophic forgetting, we combined three QA datasets-ARC-Challenge(ARC-C) (Clark et al., 2018), GSM8K (Cobbe et al., 2021), Multipl-E (Cassano et al., 2022)-as the test set , which were used training/fine-tuning LLMs we chosen. This fusion dataset will be referred to as F-task in following parts. Further details can be found in the appendix C.
LLM Backbones, LoRA variants, and Knowledge retention method.
Our study employs five state-of-art LLMs: Gemma2-2B, Qwen2-1.5B/7B, and Llama3-8B/3.1-8B.
We evaluate our SECURA framework using four PEFT variants:
  • LoRA (Whitehouse et al., 2024):Standard low-rank adaptation.
  • DoRA (Liu et al., 2024):Weight-Decomposed Low-Rank Adaptation.
  • CUR-LoRA (Fawi, 2024):CUR-Decomposed low-rank adaptation
  • I-LoRA (Ren et al., 2024):Interpolation-based low-rank adaptation.
Additionally, we implement five continuous learning baselines:
  • SEQ (Sutskever et al., 2014):Sequential Learning
  • EWC (Aich, 2021):Elastic Weight Consolidation
  • ER (Fedus et al., 2020):Experience Retention
  • CUR-LoRA (Fawi, 2024):CUR-Decomposed low-rank adaptation
  • I-LoRA (Ren et al., 2024):Interpolation-based low-rank adaptation.
Evaluation Metrics and Main Experiments
We design three main experiments to assess SECURA’s performance:
Experiment 1 (PEFT Comparison): Compares LoRA variants through 5 basic model backbones with 4 MCQ measured by accuracy (AC) and 5 QA performance measured by BLEU, ROUGE-1, ROUGE-L.
Experiment 2 (Catastrophic Forgetting with PEFT Comparison): In order to clearly evaluates knowledge retention with unified training extreme environment (learning rate 1e-3) under different PEFT method, we use the F-task set measured by BLEU, ROUGE-1, ROUGE-L.
Experiment 3 (Catastrophic Forgetting with Knowledge retention method): Benchmarks against continual learning methods at 1e-5 learning rate with 32K steps 16 task sequential training(each training 2k steps) every PEFT based training method will fusion itself to frozen basic weights(e.g. CUR-LoRA, I-LoRA, SECURA) after one training task is done, measuring remain knowledge by using the F-task set and measured by BLEU, ROUGE-1, ROUGE-L.
Note: All main experiments use the M1 merge method. For hyperparameter details, refer to the appendix D.
4.2 Main Result
Task LoRA DoRA CUR-LoRA ILoRA SECURA
MCQ Avg.
AbsNarr 83.81 85.00 64.06 80.70 87.94
ObjCount 86.49 86.43 59.35 82.49 89.40
PlayDiag 85.71 84.93 56.53 76.74 89.87
ColorReason 80.44 80.86 58.30 79.64 84.37
QA Avg.
GSM8K 81.11 81.53 66.79 38.21 82.11
NewsDE 62.06 61.81 60.68 40.60 64.20
NewsIT 65.19 65.78 63.75 42.66 66.43
NewsES 60.25 58.54 60.70 39.47 66.06
ALPACA 53.22 54.41 41.09 26.53 55.82
Table 1: Performance comparison of PEFT variants under 5 LLMs across 4 MCQ tasks and 5 QA tasks. MCQ shows accuracy, QA shows BLEU, ROUGE-1, and ROUGE-L averages across all 5 LLMs. Detailed results are in Appendix E.
Performance on PEFT Comparison
As shown in Table 1 SECURA outperforms all PEFT methods across both MCQ and QA tasks, achieving SOTA results. Compared to basic lora training, SECURA get average improvement 3.63% in MCQ tasks and gains 2.56% improvement in QA tasks. Notably, on tasks of NewsES, SECURA demonstrates nearly a 6% improvement, highlighting its exceptional fine-tuning capabilities. In contrast, other variants which major in handle the problem of catastrophic forgetting got negative improvement, even those focused on fine-tuning performance like DoRA, shows minor improvements, which only gains most improvement 1.19% in abstract narrative understanding.
Task LoRA DoRA CUR-LoRA ILoRA SECURA
AbsNarr 00.80 01.19 04.48 09.33 72.09
DisflQA 06.59 23.93 10.54 26.80 72.34
LangID 00.47 03.95 08.29 24.32 72.42
LogDeduc 01.03 00.30 19.01 25.82 72.66
ObjCount 08.10 07.88 21.44 42.59 71.18
PlayDiag 00.22 00.71 05.97 33.96 72.43
ColorReason 01.17 00.28 08.56 39.18 72.31
TrackObj 00.26 01.53 03.22 21.77 72.26
Table 2: Knowledge retention comparison with PEFT variants under 5 LLMs, using 8 tasks separated trained and tested on the F-task. Performance is measured by BLEU, ROUGE-1, and ROUGE-L averages across all 5 LLMs. Detailed results are in Appendix E.
Catastrophic Forgetting with PEFT Comparison
To clearly shows SECURA’s capability of Knowledge retention, we create a extreme environment(learning rate is 1e-3) through paper (Kirkpatrick et al., 2017)(Li and Hoiem, 2018)(Zenke et al., 2017) which shows with higher learning rate, the risk of catastrophic raised. The performance shows in the table 2.
As shown in table, most of PEFT variants almost lose all of its former knowledge, even the variants which aimed to relive catastrophic forgetting such as CUR-LoRA and I-LoRA are still have heavily forgetting and only can keep average score under 42.59%, while SECURA shows strong capabilities of limit catastrophic forgetting even in such tense environment and gains average score of 72.21% under the test of F-task. Demonstrating its robustness in mitigating catastrophic forgetting.
Catastrophic Forgetting with Knowledge retention method
SECURA is compared with other knowledge retention methods across 16 sequential training tasks, tested under the F-task. As shown in Figure 4, SECURA consistently outperforms other methods in retaining knowledge, with only slight improvements over ER and I-LoRA. However, SECURA maintains superior knowledge retention and fine-tuning capability, making it an effective method for fine-tuning LLMs while limiting catastrophic forgetting.
4.3 Ablation Study
How to Prove the merge hypothesis of the Update Cycle?
Task SECURA-1 SECURA-200 SECURA-1 SECURA-200
Qwen-2 7B Llama-3.1 8b
AbsNarr 77.93 07.36 74.73 54.17
DisflQA 78.46 70.98 73.85 66.99
LangID 78.86 00.16 73.13 4.866
LogDeduc 75.19 54.61 75.11 54.46
ObjCount 78.30 73.30 73.41 61.19
PlayDiag 78.10 00.22 74.81 62.30
ColorReason 77.81 00.26 75.10 39.13
TrackObj 78.38 74.29 74.35 65.78
Table 3: Ablation study comparing performance of 1-step and 200-step merges. The 1-step merge achieves better performance with minimal forgetting, while the 200-step merge leads to a cycle with reduced performance
To prove the cycle hypothesis, we conducted an experiment comparing the performance of 1-step and 200-step merges across different LLMs. The results are shown in the table 3. Although the 200-step merge still retains some capabilities, it has obviously lower performance compared to the 1-step merge. We hypothesize that this performance degradation occurs because in the 200-step merge, the system has started entering the second cycle but not fully tracked by the problem. This suggests that the merge function plays a critical role in preventing catastrophic forgetting by maintaining a stable learning cycle.
Different Performance based on different Merge method
Merge Under EXP1: MergeMethod-1 MergeMethod-2
AbsNarr 87.48 77.27
ObjCount 85.21 79.83
PlayDiag 90.03 74.00
ColorReason 88.53 70.37
Merge Under EXP2: MergeMethod-1 MergeMethod-2
AbsNarr 64.18 / 63.77 / 50.89 74.47 / 74.71 / 69.06
ObjCount 62.37 / 63.15 / 50.86 76.12 / 76.01 / 70.01
PlayDiag 63.56 / 64.47 / 51.53 75.15 / 75.32 / 69.32
ColorReason 64.93 / 64.64 / 51.72 75.91 / 75.90 / 69.86
Table 4: Comparison of fine-tuning efficiency and knowledge retention using MergeMethod-1 and MergeMethod-2 under two experiments (EXP1 and EXP2).
We designed another ablation study by testing different performance under M1 and M2 merge method with Gemma2 2b backbone and shows its feedback on table 4.
As the prediction in methodology, M1 method with deepen merge which will change basic weights shows more benefits on fine-tuning, performance better than M2 method range from 5.62% to 18.16%. In contrast M2 method which with shallow merge method shows more advantage under knowledge retention capabilities. Where it average improved retention 11.65% in BLEU, 11.48% in ROUGE-1 and 18.31% in ROUGE-L compared to M1 method.
5 Discussion
5.1 Why Do We Update the Basic Parameters or Track SECURA Parameters But Never Change $C$ and $R$?
We assume that parameters with higher norms carry more important information while lower-norm parameters are considered less important. The columns and rows $C$ and $R$ we initialized are represent the less important parameters’ positions. Our goal is to keep the former important knowledge intact, which means it is necessary to prevent changes to $C$ and $R$ which are used to identify the locations that should be updated. By keeping $C$ and $R$ unchanged, we ensure that the critical information in the LLM is protected, while non-critical parameters can be updated or fine-tuned as needed.
5.2 What Are the Other Benefits of SECURA Compared to Other Catastrophic Forgetting Methods?
Since we are already shows SECURA’s performance in our experiments. Here, we will point out the benefits of using SECURA over other methods like EWC or Experience Replay, specifically in terms of efficiency and simplicity.
5.2.1 Smaller Weight
Traditional methods like EWC rely on the Fisher matrix to detect changes between current and previous parameters, requiring the storage of previous parameters, which increases computational costs and space requirements. In contrast, SECURA as a PEFT fine-tuning method can be outperforming with even lower parameters compared to basic LoRA training D.
5.2.2 No Need for Former Datasets and Calculation
Methods like Experience Replay store and reuse previous datasets to prevent catastrophic forgetting, but this comes at the cost of time and computational resources. SECURA, however, does not require past data and still performs effectively. This eliminates the need for dataset storage and reuse, making SECURA a more resource-efficient solution.
5.2.3 Easy to Establish
Both EWC and Experience Replay require additional preparation steps before training begin, which adds complexity. In contrast, SECURA allows fine-tuning the model without any additional operations—just like fine-tuning a LoRA model. This makes SECURA much easier to implement and more accessible for practical applications.
6 Conclusion
We propose SECURA, a novel PEFT fine-tune method that integrates SigNorm normalization and CABR decomposition to address catastrophic forgetting in LLM fine-tuning. SECURA dynamically balances parameter updates using a Sigmoid-based pruning mechanism, preserving critical knowledge while adapting to new tasks. Experiments on 18 datasets and 5 LLMs show SECURA outperforms standard LoRA, with a 3.63% improvement on MCQ tasks and 2.56% on QA tasks, while retaining over 70% of base knowledge in continual learning scenarios. Paving the way for sustainable and ethical deployment of large AI models.
7 Limitations
  • Lack of experiments on 70B+ LLMs: Due to device limitations, further research on larger models (70B+ parameters) could not be conducted. Future work will explore scaling SECURA to more massive LLMs.
  • Computational Overhead: SigNorm introduces additional matrix operations during training, increasing per-step time by approximately 1.18% compared to LoRA. Future work could focus on optimizing the normalization layer for real-time applications.
Appendix A Broader Impacts
  • Efficiency and Accessibility: By eliminating the need for prior datasets and reducing training time for knowledge retention, SECURA lowers the computational barrier for fine-tuning LLMs, making advanced AI adaptation more accessible to researchers with limited resources.
  • Environmental Sustainability: The reduced computational overhead compared to Experience Replay could decrease energy consumption in large-scale LLM deployments.
  • Ethical Risks: While SECURA mitigates catastrophic forgetting, its ability to retain sensitive information from pre-training data (e.g., biased or private content) requires careful auditing before deployment. Future work should investigate selective forgetting mechanisms to align with ethical AI principles.
  • Generalization Beyond NLP: The SigNorm algorithm, though initially designed for PEFT, could be extended to other continual learning scenarios (e.g., robotics, healthcare), potentially improving adaptive systems in dynamic environments.
Input: Base weights $\mathcal{W}_{\text{former}}$, hyperparameters $r,m,\epsilon,\text{Scale}$
Output: Updated weights $\mathcal{W}_{\text{updated}}$
Step 1: CABR-LoRA Decomposition
1. Initialize CABR matrices:  $\mathcal{W}_{\text{frozen}}=USV^{T}$
  // SVD of base weights
$\mathcal{W}_{A}\leftarrow U[:r,:]\cdot S\cdot V[m:,:]^{T}$
  // Truncated SVD for $A$
$\mathcal{W}_{B}\leftarrow\text{Zeros}(m\times r)$
  // Zero-initialized $B$
$\mathcal{W}_{\text{CABR}}\leftarrow C\cdot\mathcal{W}_{A}\cdot\mathcal{W}_{B}\cdot R$
  // CABR reconstruction
Step 2: SigNorm Normalization
2. Merge weights:
$\mathcal{W}_{\text{SECURAMerged}}\leftarrow\mathcal{W}_{\text{CABR}}+\mathcal{W}_{\text{former}}$
3. Compute magnitude ratio:
$\mathcal{M}_{\text{Mag}}\leftarrow\left|\frac{\mathcal{W}_{\text{SECURAMerged}}}{\mathcal{W}_{\text{former}}+\epsilon}\right|$
  // Avoid division by zero
4. Normalize and scale:
$\mathcal{M}_{\text{norm}}\leftarrow\left[\frac{\mathcal{M}_{\text{Mag}}}{\max(\mathcal{M}_{\text{Mag}})+\epsilon}-0.5\right]\cdot\text{Scale}$
5. Apply Sigmoid constraint:
$\mathcal{M}_{\text{Res}}\leftarrow 2-\sigma(\mathcal{M}_{\text{norm}})$
  // $\sigma$: Sigmoid function
6. Adjust updates:
$\mathcal{W}_{\text{updated}}\leftarrow\mathcal{W}_{\text{SECURAMerged}}/\mathcal{M}_{\text{Res}}$
  // Element-wise division
Step 3: Merge Strategy (M1/M2)
7. if M1 (Deep Merge) then
        $\mathcal{W}_{\text{Newformer}}\leftarrow\mathcal{W}_{\text{CABR}}+\mathcal{W}_{\text{former}}$
        // Update base weights
end if
else if M2 (Shallow Merge) then
        $\mathcal{W}_{\text{former}}\leftarrow CA_{\text{former}}B_{\text{former}}R+\mathcal{W}_{\text{basic}}$
        $\mathcal{W}_{\text{Newformer}}\leftarrow\mathcal{W}_{\text{former}}+\mathcal{W}_{\text{CABR}}$
        // Preserve $\mathcal{W}_{\text{basic}}$
end if
return $\mathcal{W}_{\text{updated}}$
Algorithm 1 SECURA Methodology
Appendix B Pseudo code deduction
Algorithm 1 shows whole deduction of SECURA:
Appendix C Details about 18 Tasks and Datasets
Table 5 includes detailed descriptions of each dataset’s name, keywords, main content and corresponding evaluation metrics. These 18 tasks include different topics, such as mathematical reasoning, logical deduction, multilingual identification, programming challenges, translation questions, and commonsense understanding.
Task Name Keywords Description Evaluation Metrics
abstract_narrative_understanding (AbsNarr) narrative understanding, multiple choice Given a narrative, choose the most related proverb. Accuracy
cnn_dailymail (CNNDM) summarization Given news articles, write the summarization. ROUGE
disfl_qa (DisflQA) contextual question-answering, reading comprehension Pick the correct answer span from the context given the disfluent question. Accuracy
gsm8k (GSM8K) mathematics Solve the grade school math word problems. Accuracy
language_identification (LangID) multilingual, multiple choice Given a sentence, select the correct language. Accuracy
linguistics_puzzles (LingPuzz) logical reasoning, linguistics Solve Rosetta Stone-style linguistics puzzles. BLEU, ROUGE
logical_deduction (LogDeduc) logical reasoning, multiple choice Deduce the order of a sequence of objects. Accuracy
news_commentary_de (NewsDE) multilingual, translation Translate German sentences into English. BLEU
news_commentary_es (NewsES) multilingual, translation Translate Spanish sentences into English. BLEU
news_commentary_it (NewsIT) multilingual, translation Translate Italian sentences into English. BLEU
object_counting (ObjCount) logical reasoning Questions that involve enumerating objects and asking the model to count them. Accuracy
play_dialog_same_or_different (PlayDiag) reading comprehension, multiple choice Determine if nearby lines in a Shakespeare play were spoken by the same individual. Accuracy
question_selection (QuestSel) reading comprehension, multiple choice Given an answer along with its context, select the most appropriate question which has the given answer as its answer. Accuracy
reasoning_about_colored_objects (ColorReason) reading comprehension, logical reasoning, multiple choice Answer extremely simple questions about the colors of objects on a surface. Accuracy
strategyqa (StratQA) logical reasoning, context-free question answering Answer questions in which the required reasoning steps are implicit in the question. BLEU, ROUGE, Accuracy
tracking_shuffled_objects (TrackObj) logical reasoning, multiple choice Determine the final positions given initial positions and a description of a sequence of swaps. Accuracy
Abstraction and Reasoning Corpus - Challenge (ARC-C) logical reasoning, multiple choice Answer questions that require reasoning about complex, abstract scenarios, often requiring multi-step logic and understanding of concepts. Accuracy
Multiple-Choice Evaluations for Embodied Agents (Multiple-E) reasoning, multiple choice, dynamic problems Answer questions based on a sequence of logical events or transformations, with an emphasis on evolving scenarios and inference. Accuracy
Table 5: Details about the 18 selected tasks following Zhang and Li (2025).
Appendix D Experimental hyper parameters on three main experiments
Table 123,  4 and Figure 4 show all hyper-parameters among different LoRA baselines and SECURA using in three main experiments.
Hyper Parameters LoRA DoRA CUR-LoRA ILoRA SECURA
Exp 1
Rank 8 8 - 8 -
learning rates 1e-5 1e-4 1e-4 1e-3 1e-3
CABR/CUR r Rank - - 150 - 150
CABR m Rank - - - - 200
Fusion Step - - - - 1
Trainable parameters(each layer) 1.31e+5 1.48e+5 0.90e+5 2.62e+5 1.20e+5
Exp 2
Rank 8 8 - 8 -
learning rates 1e-3 1e-3 1e-3 1e-3 1e-3
CABR/CUR r Rank - - 150 - 150
CABR m Rank - - - - 200
Fusion Step - - - - 1
Trainable parameters(each layer) 1.31e+5 1.48e+5 0.90e+5 2.62e+5 1.20e+5
Exp 3
Rank 8 8 - 8 -
learning rates 1e-5 1e-5 1e-5 1e-5 1e-5
CABR/CUR r Rank - - 150 - 150
CABR m Rank - - - - 200
Fusion Step 2000 2000 2000 2000 1
Trainable parameters(each layer) 1.31e+5 1.48e+5 0.90e+5 2.62e+5 1.20e+5
Table 6: hyper-parameters of all experiments
Appendix E Experimental Results on All Datasets
Table 789,  11 and  10 show all results among different LoRA baselines and SECURA using Gemma2 2b, Qwen2 1.5B, Qwen2 7B, LLaMA3 7B and LLaMA3.1 8B backbones.
Task LoRA DoRA CUR-LoRA ILoRA SECURA LoRA DoRA CUR-LoRA ILoRA SECURA
Gemma-2 2B Llama-3 8B
AbsNarr 68.67 72.43 37.03 73.57 87.48 90.35 90.91 71.94 90.06 89.54
ObjCount 78.74 77.34 58.34 78.64 85.21 89.43 89.76 74.08 89.52 92.83
PlayDiag 72.40 72.77 53.39 37.57 90.03 91.77 90.26 73.38 88.65 92.49
ColorReason 71.00 70.93 39.27 69.90 88.53 88.07 87.11 71.43 87.09 88.60
Qwen-2 1.5B Llama-3.1 8B
AbsNarr 84.07 85.67 34.13 77.13 86.72 89.43 88.95 76.81 89.23 88.65
ObjCount 84.38 85.72 36.24 73.57 85.01 89.51 89.68 74.23 89.55 93.44
PlayDiag 85.20 84.40 40.27 83.87 85.72 91.51 91.97 71.60 89.94 92.51
ColorReason 74.60 76.47 52.60 77.22 75.38 88.12 87.46 72.07 87.14 88.34
Qwen-2 7B Avg.
AbsNarr 86.53 87.06 40.41 73.49 87.33 83.81 85.00 64.06 80.70 87.94
ObjCount 90.38 89.67 53.85 81.16 90.50 86.49 86.43 59.35 82.49 89.40
PlayDiag 87.67 85.27 44.00 83.65 88.62 85.71 84.93 56.53 76.74 89.87
ColorReason 80.40 82.33 56.13 76.83 81.00 80.44 80.86 58.30 79.64 84.37
Table 7: EXP1-MCQ
Task LoRA DoRA CUR-LoRA ILoRA SECURA
Gemma-2 2B
GSM8K 58.38 / 56.01 / 50.74 60.88 / 58.20 / 53.73 57.03 / 55.28 / 49.41 60.22 / 58.76 / 53.86 58.84 / 56.25 / 50.95
NewsDE 63.42 / 62.57 / 57.20 63.53 / 62.53 / 56.74 64.02 / 63.84 / 57.81 62.82 / 62.33 / 56.69 61.98 / 60.69 / 51.41
NewsIT 65.51 / 64.25 / 62.96 66.02 / 65.71 / 64.50 67.73 / 66.26 / 65.16 65.06 / 63.68 / 62.25 61.94 / 61.10 / 59.42
NewsES 58.00 / 55.58 / 51.52 58.40 / 55.98 / 51.99 61.23 / 61.53 / 57.01 61.79 / 61.58 / 57.34 58.78 / 55.85 / 50.81
ALPACA 40.12 / 25.18 / 22.50 39.62 / 26.77 / 24.14 41.94 / 27.71 / 22.35 42.77 / 28.13 / 25.21 38.92 / 23.54 / 20.10
Qwen-2 1.5B
GSM8K 84.23 / 86.60 / 85.15 84.08 / 86.48 / 84.98 79.93 / 84.02 / 81.54 7.82 / 14.80 / 12.37 86.62 / 87.73 / 85.15
NewsDE 64.39 / 63.63 / 62.16 64.72 / 64.02 / 62.82 60.77 / 58.30 / 57.15 5.09 / 11.68 / 11.07 63.31 / 64.62 / 64.63
NewsIT 65.34 / 65.34 / 63.58 66.54 / 66.74 / 64.90 64.37 / 63.15 / 61.27 4.38 / 9.35 / 8.57 65.98 / 65.43 / 64.27
NewsES 65.74 / 66.71 / 65.57 64.82 / 65.19 / 64.05 61.97/ 61.39 / 49.41 5.67 / 14.39 / 12.36 65.88 / 67.31 / 66.74
ALPACA 65.49 / 61.63 / 55.08 65.78 / 62.00 / 55.50 57.03 / 55.28 / 49.41 4.80 / 15.53 / 12.80 66.82 / 64.31 / 55.19
Qwen-2 7B
GSM8K 91.94 / 94.54 / 94.37 91.80 / 94.27 / 94.17 57.03 / 55.28 / 49.41 2.40 / 5.36 / 4.68 94.33 / 94.98 / 95.00
NewsDE 66.39 / 66.15 / 64.72 66.87 / 66.86 / 65.71 57.03 / 55.28 / 49.41 6.59 / 8.68 / 7.88 68.83 / 67.70 / 64.74
NewsIT 68.72 / 69.00 / 67.81 69.17 / 69.32 / 68.23 57.03 / 55.28 / 49.41 6.24 / 9.04 / 8.90 68.54 / 70.30 / 69.34
NewsES 67.96 / 69.49 / 68.80 67.37 / 68.77 / 68.08 57.03 / 55.28 / 49.41 7.26 / 10.21 / 9.73 68.89 / 69.94 / 70.00
ALPACA 66.55 / 63.14 / 57.42 67.17/ 64.05 / 58.38 57.03 / 55.28 / 49.41 6.61 / 12.78 / 11.32 68.25 / 65.64 / 58.86
Llama-3 8B
GSM8K 84.53 / 87.62 / 85.96 83.21 / 88.20 / 86.73 59.73 / 56.81 / 52.22 61.92 / 59.31 / 55.49 86.67 / 88.72 / 86.54
NewsDE 63.32 / 61.20 / 56.48 61.88 / 60.29 / 55.33 63.50 / 61.69 / 56.32 65.34 / 63.58 / 58.48 66.10 / 65.68 / 64.60
NewsIT 65.12 / 64.00 / 63.03 64.92 / 64.37 / 63.45 66.01 / 64.52 / 63.49 67.11 / 67.23 / 66.49 68.63/ 69.08 / 67.39
NewsES 59.15 / 55.98 / 52.04 53.69 / 53.25 / 49.24 62.48 / 60.84 / 56.36 60.81 / 58.45 / 54.89 68.97 / 69.97 / 69.20
ALPACA 62.11 / 57.03 / 51.20 64.15 / 58.91 / 55.22 37.11 / 23.24 / 19.95 48.06 / 36.66 / 34.83 65.84 / 62.83 / 57.11
Llama-3.1 8B
GSM8K 84.41 / 87.29 / 85.02 84.25 / 86.89 / 85.11 60.36 / 57.60 / 53.31 62.09 / 58.99 / 55.20 86.03 / 87.93 / 86.04
NewsDE 62.85 / 60.54 / 55.85 61.19 / 59.72 / 54.93 65.80 / 62.57 / 57.31 65.93 / 63.98 / 58.92 66.02 / 66.86 / 65.87
NewsIT 65.35 / 64.50 / 63.40 64.98 / 64.33 / 63.47 65.12 / 63.73 / 62.75 67.69 / 67.40 / 66.65 68.27 / 69.18 / 67.63
NewsES 59.31 / 55.94 / 52.08 54.25 / 53.39 / 49.73 61.36 / 60.00 / 55.90 62.10 / 59.63 / 55.87 69.05 / 70.29 / 69.25
ALPACA 61.03 / 58.92 / 51.09 61.93 / 60.21 / 52.33 37.89 / 23.94 / 20.89 48.67 / 35.96 / 33.87 67.10 / 64.11 / 58.77
Avg.
GSM8K 80.69 / 82.41 / 80.24 80.84 / 82.80 / 80.94 67.84 / 68.12 / 64.42 38.89 / 39.44 / 36.32 82.49 / 83.12 / 80.73
NewsDE 64.08 / 62.82 / 59.28 63.64 / 62.69 / 59.11 63.07 / 61.41 / 57.56 41.16 / 42.05 / 38.61 65.25 / 65.11 / 62.25
NewsIT 66.01 / 65.42 / 64.16 66.33 / 66.09 / 64.93 65.03/ 63.77 / 62.47 42.10 / 43.34 / 42.57 66.67 / 67.02 / 65.61
NewsES 62.03 / 60.74 / 58.00 59.70 / 59.31 / 56.62 62.15 / 61.49 / 58.48 39.53 / 40.85 / 38.04 66.31 / 66.67 / 65.20
ALPACA 59.05 / 53.17 / 47.45 59.73 / 54.38 / 49.11 49.13 / 39.55 / 34.61 30.18 / 25.81 / 23.61 61.39 / 56.09 / 50.00
Table 8: EXP1-QA BLEU/Rouge1/RougeL
Task LoRA DoRA CUR-LoRA ILoRA SECURA
Gemma-2 2B
AbsNarr 00.09 / 00.24 / 00.24 00.17 / 00.11 / 00.11 19.91 / 20.13 / 16.41 46.01 / 49.41 / 42.03 64.18 / 63.77 / 50.89
DisflQA 00.10 / 00.58 / 00.57 00.19 / 00.52 / 00.50 08.78 / 11.92 / 09.44 65.92 / 66.79 / 60.03 65.16 / 65.00 / 52.01
LangID 00.30 / 00.52 / 00.50 00.06 / 00.20 / 00.20 31.57 / 39.89 / 32.39 66.61 / 67.49 / 60.81 65.16 / 64.88 / 52.19
LogDeduc 03.14 / 04.25 / 03.71 02.01 / 00.26 / 00.25 17.23 / 22.95 / 17.70 68.95 / 70.12 / 63.87 66.86 / 67.70 / 55.27
ObjCount 30.59 / 30.54 / 21.41 39.57 / 42.87 / 32.58 49.18 / 52.11 / 40.85 68.59 / 69.04 / 62.03 62.37 / 63.15 / 50.86
PlayDiag 00.07 / 01.23 / 01.21 00.15 / 01.73 / 01.63 16.33 / 19.69 / 14.99 60.30 / 62.69 / 56.41 63.56 / 64.47 / 51.53
ColorReason 02.96 / 06.16 / 05.10 00.59 / 01.36 / 01.26 23.94 / 30.21 / 22.94 62.85 / 63.64 / 57.69 64.93 / 64.64 / 51.72
TrackObj 00.12 / 01.19 / 01.17 00.16 / 00.38 / 00.36 01.10 / 03.73 / 03.34 41.25 / 42.73 / 35.41 64.28 / 64.67 / 51.79
Qwen-2 1.5B
AbsNarr 03.75 / 03.72 / 03.31 06.35 / 05.69 / 04.82 00.28 / 00.11 / 00.11 00.82 / 00.39 / 00.36 75.07 / 75.83 / 70.80
DisflQA 29.04 / 37.23 / 29.42 47.51 / 48.82 / 40.14 34.54 / 38.93 / 30.47 00.32 / 01.03 / 01.00 75.83 / 76.99 / 72.15
LangID 00.50 / 00.00 / 00.00 00.26 / 00.00 / 00.00 00.24 / 00.19 / 00.19 00.60 / 01.70 / 01.63 76.75 / 76.98 / 72.44
LogDeduc 00.26 / 00.34 / 00.34 00.29 / 00.19 / 00.19 00.27 / 00.87 / 00.80 01.29 / 01.75 / 01.52 76.92 / 77.21 / 71.87
ObjCount 01.12 / 00.14 / 00.14 00.84 / 00.00 / 00.00 06.99 / 08.61 / 07.43 02.43 / 04.65 / 03.77 73.31 / 75.28 / 70.46
PlayDiag 00.18 / 00.00 / 00.00 00.31 / 00.00 / 00.00 00.27 / 00.01 / 00.01 00.29 / 00.72 / 00.70 74.98 / 76.57 / 72.04
ColorReason 00.35 / 00.00 / 00.00 00.25 / 00.04 / 00.04 04.81 / 08.01 / 06.27 00.29 / 00.02 / 00.02 75.01 / 75.78 / 70.83
TrackObj 00.62 / 00.00 / 00.00 05.11 / 08.39 / 07.84 00.42 / 00.72 / 00.66 00.48 / 00.62 / 00.60 74.71 / 76.28 / 71.37
Qwen-2 7B
AbsNarr 00.28 / 00.05 / 00.05 00.22 / 00.12 / 00.12 00.67 / 00.35 / 00.35 00.18 / 00.00 / 00.00 78.18 / 79.74 / 75.87
DisflQA 00.29 / 00.16 / 00.16 41.88 / 44.69 / 34.05 01.00 / 00.59 / 00.54 00.28 / 00.00 / 00.00 78.46 / 80.25 / 76.70
LangID 00.68 / 00.00 / 00.00 00.65 / 00.00 / 00.00 00.29 / 00.01 / 00.01 00.68 / 00.00 / 00.00 78.58 / 80.88 / 77.13
LogDeduc 00.50 / 00.85 / 00.85 00.30 / 00.14 / 00.14 49.70 / 54.73 / 49.87 00.52 / 00.22 / 00.22 75.41 / 77.01 / 73.15
ObjCount 00.95 / 00.00 / 00.00 01.32 / 00.10 / 00.10 00.68 / 00.05 / 00.05 00.88 / 00.00 / 00.00 77.71 / 80.64 / 76.56
PlayDiag 00.21 / 00.00 / 00.00 00.63 / 00.00 / 00.00 00.55 / 00.56 / 00.56 00.30 / 00.03 / 00.03 77.39 / 80.31 / 76.62
ColorReason 00.65 / 00.02 / 00.02 00.20 / 00.00 / 00.00 00.30 / 00.00 / 00.00 00.22 / 00.02 / 00.02 78.09 / 79.62 / 75.72
TrackObj 00.38 / 00.02 / 00.02 00.01 / 00.00 / 00.00 08.70 / 07.30 / 06.39 00.27 / 00.09 / 00.09 78.06 / 80.36 / 76.74
Llama-3 8B
AbsNarr 00.15 / 00.06 / 00.06 00.07 / 00.01 / 00.01 01.90 / 03.47 / 03.13 00.08 / 00.30 / 00.30 75.32 / 76.84 / 70.77
DisflQA 00.08 / 00.15 / 00.15 00.05 / 00.44 / 00.44 00.89 / 04.11 / 03.93 00.11 / 01.78 / 01.71 74.32 / 76.77 / 69.94
LangID 00.14 / 00.03 / 00.03 16.73 / 21.59 / 18.88 01.38 / 07.69 / 07.23 00.08 / 00.01 / 00.01 74.98 / 76.66 / 70.29
LogDeduc 00.88 / 00.07 / 00.07 00.52 / 00.06 / 00.06 10.75 / 12.65 / 09.71 59.44 / 62.69 / 55.91 75.30 / 76.90 / 71.00
ObjCount 09.13 / 06.23 / 05.57 00.51 / 00.03 / 00.03 34.71 / 38.38 / 26.96 74.02 / 75.15 / 68.88 73.36 / 74.92 / 68.97
PlayDiag 00.07 / 00.05 / 00.05 00.20 / 02.25 / 02.25 03.50 / 08.47 / 07.40 62.97 / 65.34 / 57.40 75.31 / 77.47 / 71.76
ColorReason 00.19 / 01.09 / 00.97 00.11 / 00.17 / 00.17 06.65 / 10.82 / 08.78 66.22 / 68.48 / 59.60 75.79 / 76.93 / 70.36
TrackObj 00.20 / 00.02 / 00.02 00.13 / 00.08 / 00.07 00.08 / 00.18 / 00.18 69.13 / 71.03 / 63.49 74.80 / 76.94 / 70.88
Llama-3.1 8B
AbsNarr 00.01 / 00.00 / 00.00 00.10 / 00.00 / 00.00 00.22 / 00.04 / 00.04 00.07 / 00.03 / 00.03 75.61 / 77.14 / 71.44
DisflQA 00.05 / 00.46 / 00.46 32.60 / 38.93 / 28.23 01.69 / 05.85 / 05.58 68.98 / 70.84 / 63.32 74.45 / 77.02 / 70.10
LangID 00.09 / 02.16 / 02.14 00.10 / 00.34 / 00.34 00.25 / 01.55 / 01.55 56.69 / 58.30 / 50.32 74.43 / 75.85 / 69.13
LogDeduc 00.18 / 00.05 / 00.05 00.09 / 00.00 / 00.00 11.53 / 14.63 / 11.81 00.67 / 00.08 / 00.08 75.84 / 77.74 / 71.77
ObjCount 06.98 / 04.58 / 04.20 00.29 / 00.00 / 00.00 18.15 / 21.50 / 16.06 70.13 / 72.79 / 66.51 74.22 / 76.07 / 69.96
PlayDiag 00.07 / 00.09 / 00.09 00.21 / 00.67 / 00.67 02.33 / 07.99 / 06.87 47.87 / 51.97 / 42.48 75.13 / 77.56 / 71.74
ColorReason 00.05 / 00.03 / 00.03 00.10 / 00.00 / 00.00 00.75 / 02.58 / 02.30 71.29 / 72.11 / 65.31 76.65 / 77.63 / 71.04
TrackObj 00.04 / 00.09 / 00.09 00.10 / 00.15 / 00.15 01.06 / 07.59 / 06.86 00.16 / 00.62 / 00.59 74.90 / 77.09 / 71.06
Avg.
AbsNarr 00.85 / 00.81 / 00.73 01.38 / 01.19 / 01.01 04.60 / 04.82 / 04.01 09.43 / 10.03 / 08.54 73.67 / 74.66 / 67.95
DisflQA 05.91 / 07.72 / 06.15 24.44 / 26.68 / 20.67 09.38 / 12.28 / 09.99 27.12 / 28.09 / 25.21 73.65 / 75.20 / 68.18
LangID 00.34 / 00.54 / 00.53 03.56 / 04.42 / 03.88 06.74 / 09.86 / 08.27 24.93 / 25.50 / 22.55 73.98 / 75.05 / 68.24
LogDeduc 00.99 / 01.11 / 01.00 00.64 / 00.13 / 00.13 17.89 / 21.17 / 17.98 26.18 / 26.97 / 24.32 74.07 / 75.31 / 68.61
ObjCount 09.75 / 08.30 / 06.26 08.51 / 08.60 / 06.54 21.94 / 24.13 / 18.27 43.21 / 44.32 / 40.24 72.19 / 74.01 / 67.36
PlayDiag 00.12 / 00.27 / 00.27 00.30 / 00.93 / 00.91 04.59 / 07.35 / 05.96 34.35 / 36.15 / 31.40 73.27 / 75.28 / 68.74
ColorReason 00.84 / 01.46 / 01.22 00.25 / 00.31 / 00.29 07.29 / 10.32 / 08.06 40.17 / 40.85 / 36.53 74.09 / 74.92 / 67.93
TrackObj 00.27 / 00.26 / 00.26 01.10 / 01.80 / 01.68 02.27 / 03.90 / 03.48 22.26 / 23.02 / 20.04 73.35 / 75.07 / 68.37
Table 9: EXP2-Forgetting BLEU/Rouge1/RougeL
Datasets SEQ EWC ER I-LORA SECURA
AbsNarr-1k 73.28 / 75.55 / 68.58 75.39 / 58.84 / 59.90 67.77 / 75.14 / 69.68 75.57 / 76.75 / 70.81 78.40 / 79.69 / 75.94
AbsNarr-2k 66.26 / 64.33 / 61.01 70.19 / 58.05 / 55.68 76.42 / 77.77 / 70.94 75.25 / 76.69 / 71.20 78.01 / 79.28 / 75.86
DisflQA-1k 64.90 / 65.93 / 60.63 67.96 / 67.08 / 60.41 72.77 / 75.70 / 73.73 74.96 / 75.95 / 70.25 77.96 / 78.94 / 74.97
DisflQA-2k 54.93 / 47.38 / 42.97 62.43 / 68.62 / 67.01 71.68 / 73.71 / 70.32 74.93 / 76.06 / 70.26 76.46 / 77.96 / 73.85
CNNDM-1k 46.59 / 48.36 / 37.78 68.43 / 62.65 / 62.91 69.43 / 77.45 / 70.41 74.92 / 75.74 / 69.97 78.55 / 79.21 / 75.39
CNNDM-2k 37.53 / 31.12 / 24.50 67.45 / 58.24 / 64.16 73.04 / 69.97 / 73.83 74.35 / 75.85 / 70.13 78.29 / 78.87 / 74.77
LangID-1k 28.68 / 24.94 / 25.06 68.48 / 59.60 / 66.68 78.57 / 78.31 / 72.06 74.29 / 75.70 / 69.92 76.43 / 77.76 / 73.87
LangID-2k 25.54 / 18.47 / 19.13 74.32 / 64.99 / 55.31 74.75 / 74.98 / 70.12 74.97 / 76.12 / 70.23 77.05 / 78.35 / 74.55
LingPuzz-1k 21.64 / 15.06 / 13.97 62.64 / 55.22 / 59.36 72.33 / 69.01 / 68.93 70.89 / 71.89 / 66.51 77.44 / 78.50 / 74.29
LingPuzz-2k 17.24 / 14.78 / 14.79 70.94 / 56.87 / 59.37 69.88 / 77.20 / 67.92 71.32 / 72.19 / 66.55 77.84 / 79.69 / 75.65
LogDeduc-1k 12.28 / 21.61 / 10.62 74.88 / 68.39 / 55.98 68.73 / 75.06 / 69.96 67.38 / 69.78 / 64.14 76.11 / 77.66 / 73.49
LogDeduc-2k 12.18 / 15.29 / 11.70 64.50 / 56.40 / 65.20 74.25 / 72.63 / 65.04 68.26 / 69.99 / 64.19 75.81 / 77.55 / 73.58
NewsDE-1k 18.56 / 19.40 / 10.89 66.71 / 72.97 / 53.78 75.47 / 76.34 / 71.89 69.49 / 70.59 / 64.94 77.17 / 78.14 / 73.93
NewsDE-2k 12.48 / 16.91 / 15.00 73.97 / 55.39 / 61.35 73.69 / 72.29 / 66.30 70.12 / 70.68 / 64.76 77.29 / 78.04 / 73.90
NewsES-1k 13.14 / 13.92 / 09.41 76.47 / 65.25 / 66.89 70.59 / 76.85 / 70.16 71.13 / 72.11 / 66.38 78.52 / 79.12 / 75.09
NewsES-2k 17.57 / 18.70 / 15.63 74.44 / 71.62 / 56.97 75.49 / 77.96 / 71.54 70.37 / 71.27 / 65.09 77.32 / 78.13 / 74.06
NewsIT-1k 16.79 / 15.78 / 15.28 70.15 / 60.99 / 65.53 73.05 / 69.87 / 69.61 69.73 / 71.28 / 64.91 77.04 / 77.81 / 73.60
NewsIT-2k 16.21 / 13.61 / 14.03 59.92 / 64.77 / 56.84 68.92 / 74.91 / 70.74 69.79 / 70.75 / 64.27 77.19 / 78.18 / 74.08
ObjCount-1k 20.04 / 19.88 / 12.88 58.89 / 54.79 / 59.17 73.25 / 75.17 / 73.22 70.01 / 71.12 / 64.56 75.87 / 76.80 / 72.09
ObjCount-2k 16.21 / 14.24 / 10.50 61.86 / 70.82 / 53.73 76.96 / 74.24 / 70.42 69.75 / 71.01 / 64.33 75.23 / 76.19 / 71.88
PlayDiag-1k 20.20 / 20.68 / 14.39 66.16 / 57.02 / 61.08 76.24 / 74.59 / 73.92 69.94 / 71.24 / 64.15 73.15 / 73.39 / 68.69
PlayDiag-2k 16.54 / 14.07 / 09.09 68.42 / 67.72 / 66.00 71.71 / 76.02 / 69.82 70.16 / 71.32 / 64.23 73.10 / 74.33 / 69.90
QuesSel-1k 16.48 / 18.08 / 08.91 76.35 / 64.25 / 56.73 72.41 / 78.06 / 71.14 66.44 / 67.81 / 61.11 74.90 / 76.64 / 72.43
QuesSel-2k 11.53 / 13.89 / 13.83 71.47 / 59.96 / 57.23 77.98 / 71.05 / 68.51 66.31 / 67.98 / 61.34 75.19 / 76.92 / 72.87
ColorReason-1k 20.86 / 20.52 / 08.99 62.26 / 60.83 / 56.48 68.15 / 70.82 / 75.76 61.76 / 64.05 / 57.77 76.84 / 77.60 / 73.47
ColorReason-2k 19.30 / 21.79 / 12.51 68.17 / 63.93 / 55.37 75.77 / 71.09 / 68.59 61.77 / 63.61 / 57.11 76.59 / 77.66 / 73.61
StratQA-1k 14.51 / 20.64 / 09.89 73.60 / 55.71 / 59.90 73.58 / 74.30 / 65.52 69.63 / 70.62 / 63.77 78.36 / 79.07 / 74.86
StratQA-2k 16.54 / 16.59 / 16.79 64.61 / 57.61 / 66.30 77.37 / 71.37 / 72.18 71.34 / 71.79 / 64.86 79.14 / 79.83 / 75.58
TrackObj-1k 18.87 / 18.21 / 13.45 74.40 / 64.98 / 61.91 67.89 / 74.82 / 70.75 69.08 / 70.41 / 63.87 76.03 / 77.06 / 72.91
TrackObj-2k 19.24 / 17.74 / 16.63 68.89 / 65.52 / 61.10 72.40 / 73.86 / 69.43 69.60 / 70.37 / 63.30 75.55 / 76.38 / 72.32
ALPACA-1k 19.74 / 14.36 / 15.40 67.77 / 63.03 / 62.00 73.81 / 74.09 / 67.03 69.07 / 70.09 / 62.71 78.32 / 78.97 / 74.92
ALPACA-2k 18.90 / 17.15 / 15.64 64.89 / 62.35 / 61.67 70.38 / 69.34 / 69.43 69.87 / 70.60 / 63.67 78.53 / 79.79 / 75.93
Table 10: EXP3 Qwen-2 7B
Datasets SEQ EWC ER I-LORA SECURA
AbsNarr-1k 69.18 / 63.32 / 54.88 65.54 / 61.29 / 58.09 70.07 / 73.84 / 74.78 75.57 / 76.75 / 70.81 76.95 / 78.11 / 71.80
AbsNarr-2k 59.20 / 47.47 / 55.10 62.06 / 58.39 / 62.74 67.46 / 71.10 / 72.56 75.25 / 76.69 / 71.20 75.70 / 76.96 / 70.62
DisflQA-1k 55.71 / 46.66 / 47.68 63.52 / 58.81 / 58.65 68.45 / 75.35 / 68.83 74.96 / 75.95 / 70.25 75.58 / 76.86 / 70.58
DisflQA-2k 54.12 / 41.02 / 51.27 60.74 / 59.32 / 63.52 73.34 / 74.68 / 67.86 74.93 / 76.06 / 70.26 74.83 / 76.61 / 70.34
CNNDM-1k 49.91 / 40.61 / 45.57 68.43 / 63.42 / 63.54 73.78 / 75.84 / 67.71 74.92 / 75.74 / 69.97 76.36 / 78.03 / 71.65
CNNDM-2k 45.65 / 34.81 / 43.70 67.03 / 60.14 / 57.00 73.42 / 74.05 / 70.98 74.35 / 75.85 / 70.13 76.06 / 77.52 / 71.24
LangID-1k 41.22 / 33.40 / 28.07 64.08 / 63.61 / 58.88 73.27 / 74.73 / 69.55 74.29 / 75.70 / 69.92 73.83 / 76.12 / 69.80
LangID-2k 37.41 / 33.45 / 26.26 68.13 / 58.36 / 57.56 74.64 / 79.49 / 66.13 74.97 / 76.12 / 70.23 73.55 / 75.70 / 69.52
LingPuzz-1k 37.15 / 29.39 / 30.99 60.18 / 67.05 / 56.20 70.67 / 78.64 / 69.14 70.89 / 71.89 / 66.51 75.77 / 76.93 / 70.69
LingPuzz-2k 28.57 / 24.65 / 24.37 59.40 / 58.79 / 56.76 72.65 / 73.37 / 71.63 71.32 / 72.19 / 66.55 75.43 / 76.46 / 70.42
LogDeduc-1k 26.00 / 13.26 / 16.88 67.22 / 58.51 / 58.65 76.15 / 78.50 / 68.74 67.38 / 69.78 / 64.14 75.48 / 76.63 / 70.70
LogDeduc-2k 19.16 / 07.08 / 16.08 64.65 / 60.56 / 56.30 71.59 / 71.41 / 70.79 68.26 / 69.99 / 64.19 76.25 / 77.13 / 70.94
NewsDE-1k 14.37 / 10.09 / 15.45 64.15 / 60.38 / 60.14 73.55 / 75.26 / 70.54 69.49 / 70.59 / 64.94 75.92 / 76.90 / 70.82
NewsDE-2k 15.66 / 06.82 / 13.88 59.85 / 61.09 / 57.85 68.35 / 79.48 / 66.44 70.12 / 70.68 / 64.76 76.35 / 77.32 / 71.30
NewsES-1k 14.11 / 12.49 / 06.47 59.54 / 64.15 / 57.66 70.96 / 73.53 / 69.15 71.13 / 72.11 / 66.38 76.92 / 78.02 / 71.96
NewsES-2k 17.65 / 09.82 / 16.77 67.33 / 64.66 / 55.14 68.49 / 70.50 / 65.03 70.37 / 71.27 / 65.09 75.87 / 77.30 / 71.52
NewsIT-1k 12.58 / 09.18 / 17.16 68.69 / 64.00 / 62.40 78.12 / 75.56 / 68.42 69.73 / 71.28 / 64.91 76.25 / 77.36 / 71.19
NewsIT-2k 10.55 / 05.95 / 08.20 69.35 / 60.45 / 62.54 68.89 / 74.56 / 69.71 69.79 / 70.75 / 64.27 76.37 / 77.48 / 71.27
ObjCount-1k 16.21 / 06.25 / 09.35 58.48 / 67.04 / 54.52 72.91 / 72.87 / 68.49 70.01 / 71.12 / 64.56 75.84 / 77.38 / 71.64
ObjCount-2k 15.88 / 07.88 / 08.46 62.76 / 60.67 / 57.22 72.70 / 74.70 / 73.83 69.75 / 71.01 / 64.33 74.11 / 76.30 / 70.45
PlayDiag-1k 18.61 / 11.20 / 07.12 65.28 / 64.65 / 55.17 72.49 / 79.93 / 67.73 69.94 / 71.24 / 64.15 75.47 / 76.68 / 70.63
PlayDiag-2k 12.40 / 06.14 / 15.20 67.52 / 60.96 / 57.17 72.43 / 75.20 / 65.62 70.16 / 71.32 / 64.23 74.67 / 76.26 / 70.04
QuesSel-1k 12.86 / 12.78 / 08.42 68.95 / 57.29 / 56.37 68.78 / 70.81 / 65.90 66.44 / 67.81 / 61.11 75.10 / 76.61 / 70.44
QuesSel-2k 15.07 / 06.14 / 12.65 69.17 / 60.35 / 64.00 68.59 / 78.57 / 75.98 66.31 / 67.98 / 61.34 75.08 / 76.64 / 70.40
ColorReason-1k 13.96 / 09.85 / 09.35 59.17 / 64.56 / 57.21 72.10 / 69.43 / 73.29 61.76 / 64.05 / 57.77 74.60 / 76.52 / 70.37
ColorReason-2k 13.76 / 11.97 / 11.27 60.30 / 58.24 / 58.42 73.58 / 71.75 / 76.00 61.77 / 63.61 / 57.11 75.05 / 76.47 / 70.53
StratQA-1k 16.80 / 06.01 / 09.41 67.17 / 65.51 / 66.36 72.31 / 78.95 / 67.90 69.63 / 70.62 / 63.77 75.76 / 77.33 / 71.64
StratQA-2k 15.74 / 10.51 / 15.28 61.96 / 63.57 / 53.82 71.80 / 73.93 / 66.00 71.34 / 71.79 / 64.86 76.32 / 77.51 / 71.45
TrackObj-1k 17.79 / 14.76 / 08.23 70.91 / 57.14 / 56.88 73.40 / 77.29 / 73.51 69.08 / 70.41 / 63.87 76.99 / 77.77 / 71.78
TrackObj-2k 12.14 / 08.31 / 17.06 58.71 / 55.59 / 64.89 69.81 / 70.79 / 70.75 69.60 / 70.37 / 63.30 77.38 / 78.09 / 72.19
ALPACA-1k 12.90 / 09.89 / 06.41 66.69 / 60.25 / 63.42 74.06 / 76.02 / 69.19 69.07 / 70.09 / 62.71 76.96 / 77.55 / 71.54
ALPACA-2k 13.25 / 15.62 / 12.66 66.88 / 59.11 / 66.27 68.09 / 80.43 / 73.90 69.87 / 70.60 / 63.67 75.12 / 76.38 / 70.43
Table 11: EXP3 Llama3 8B